Selected financial data for Quick Sell, Inc., a retail store, appear as follows. Year 2 Year 1 Sales (all on account) $ 795,000 $ 601,000 Cost of goods sold 484,000 419,000 Average inventory during the year 135,000 125,000 Average receivables during the year 150,000 100,000

a-1. Compute the gross profit percentage for both years. (Round your percentage answers to the nearest whole number. i.e. 0.1234 as 12%.)

a-2. Compute the inventory turnover for both years. (Round your answers to 1 decimal place.)

a-3. Compute the accounts receivable turnover for both years. (Round your answers to 1 decimal place.)

b. Which of the following show a positive or negative trend?
Answer #1
a-1) Calculation of gross profitpercentage:
Gross profit percentage= gross profit/sales*100
Year 1
Gross profit= sales- cost of goods sold
Grss profit= 601000-419000= 182000
Gross profit percentage= 182000/601000*100= 30%
Year 2
Gross profit= sales- cost of goods sold
Gross profit= 795000-484000= 311000
Gross profit percentage= 311000/795000*100= 39%
a-2)Calculation of Inventory turnover:
inventory turnover= cost of goods sold/ Average account inventory
Year 1
inventory turnover= 419000/125000= 3.4
Year 2
inventory turnover= 484000/135000= 3.6
a-3)Calculation of Recievables turnover:
Recievables turnover= Net sales/ Average account recievables
Year 1
Recievables turnover= 601000/100000= 6
Year 2
Recievables turnover= 795000/150000= 5.3
b) Gross profit percentage and inventory turnover shows a positive trend.
Receivable turnover show negative trend.